OpenStudy (ilovepurple42043):

i cant figure out the answers to these questions ive been trying to solve them since yesterday morning. someone please help!!?? 1. write the polyomial in sandard fom. Then name the polyomial based on its degree ad number of terms. 6 - 12x + 13x^2 - 4x^2 A. 9x^2 - 12x; quadratic binomial B. 9x^2 - 12x + 6; qadratic trinomial C. -3x^2 + 6; quadratic binomial D. 9x^2 - 12x - 6; cubic trinomial 2. A bioligist studied the populations of common guppies and endlers guppies over a 6-year period. The biologist modeled the populations, in tens of thousands, with the following polynomi

OpenStudy (doc.brown):

If you have \(x\times x\) you can say \(x^2\), right?

OpenStudy (ilovepurple42043):

yea

OpenStudy (doc.brown):

Or \(x\times x\times x\times x\times x=x^5\) We use the little number to say how many times we multiply the same number over and over.

OpenStudy (ilovepurple42043):

ok

OpenStudy (ilovepurple42043):

so is it lke 29x^5 or im not sure

OpenStudy (doc.brown):

If we do that three times we can cay\[3x^2=(x\times x)+(x\times x)+(x\times x)\]

OpenStudy (ilovepurple42043):

oh okay

OpenStudy (doc.brown):

or \(3x^2+2x^2 = (x\times x)+(x\times x)+(x\times x)+(x\times x)+(x\times x)\) which is just five times doing that \((x\times x)\)

OpenStudy (ilovepurple42043):

ok that makes a whole lot more since now

OpenStudy (doc.brown):

You have \(6 - 12x + 13x^2 - 4x^2\) and see you have 13-4 of the x squared

OpenStudy (ilovepurple42043):

so 13x^2 is like saying (x times x) 13 times

OpenStudy (doc.brown):

yup

OpenStudy (ilovepurple42043):

so now its 9x^2

OpenStudy (ilovepurple42043):

?

OpenStudy (doc.brown):

Quite literally 9 times (x times x). Yes

OpenStudy (ilovepurple42043):

ok thanks!! (:

OpenStudy (doc.brown):

Do you own a bicycle?

OpenStudy (ilovepurple42043):

yea y?

OpenStudy (doc.brown):

What about a tricycle?

OpenStudy (ilovepurple42043):

uh no

OpenStudy (doc.brown):

So you only own one polycycle?

OpenStudy (ilovepurple42043):

polycycle?

OpenStudy (doc.brown):

Poly, some number of. Like polynomial, here you have three different parts, so you have a tri-nomial.

OpenStudy (ilovepurple42043):

oh ok

OpenStudy (ilovepurple42043):

wat about cubic or quadratic? how do i know that?

OpenStudy (doc.brown):

quadratus is latin for square

OpenStudy (ilovepurple42043):

ok

OpenStudy (doc.brown):

cubic means cubed or to the three \(x^3\)

OpenStudy (ilovepurple42043):

ok i forgot about that thank you so so so so much!!

OpenStudy (doc.brown):

Happy to help